# Llama-Nemotron: Efficient Reasoning Models

### v1

@@ 2.1 Deployment Constraints and Efficiency Targets
Figure 4 illustrates the trade-off between GPQA-Diamond accuracy (%) and processing throughput (tokens/s) for under two settings. Notably, LN-Ultra consistently outperforms the DeepSeek-R1 and Llama-3.1-405B in both accuracy and efficiency across these settings, clearly positioning it as a superior choice on the accuracy-throughput Pareto curve.
@@ 4.2 Model-Specific Training
differently from other models below, undergoes a three-stage SFT pipeline with a global batch size of 256 using sequence packing with effective sequence length of 32k tokens. In the first stage, the model is fine-tuned exclusively on reasoning data from code, math, and science domains (Section 3.1) with a learning rate of $1\mathrm{e}{-4}$ for four epochs. This prevents failure modes such as repetitive completions. In the second stage, we introduce non-reasoning data (Section 3.2) mixed with reasoning samples, allowing the model to learn reasoning control. In the final stage, a smaller blend focused on chat, instruction-following, and tool-calling is used.
@@
is trained on the full SFT dataset for a single epoch using a fixed learning rate of $5\mathrm{e}{-6}$, sequence length of 16k and a global batch size of 256. Smaller-scale runs suggest that performance improves up to 3–4 epochs with larger learning rates ($5\mathrm{e}{-5}$), but training was constrained by computational and time limits. Recent works (Wen et al., 2025) show that rejection fine-tuning can further improve performance; however, it does not yield gains in our experiments and is therefore omitted.
@@
is trained on the full dataset using sequence packing with effective sequence length of 24k and a global batch size of 256 to maximize token throughput—an essential strategy when fine-tuning large models with long-context reasoning data. Initial ablation runs indicated that higher learning rates such as $5\mathrm{e}{-5}$ generally improve outcomes, but consistently high learning rates caused instability, including gradient explosions. To mitigate this, we implement a linear warmup to $1\mathrm{e}{-5}$, followed by cosine decay to $1\mathrm{e}{-6}$ with a warmup ratio of 10%. Despite these measures, training encountered gradient explosions and numerical instability after the first epoch. This required training resumption with reinitialized optimizer states, after which successful convergence was achieved.
@@ 5 RL for Reasoning
As described in Section 4 and illustrated in Table 5, models can acquire strong capabilities through supervised fine-tuning by distilling knowledge from powerful teachers. However, distillation inherently sets an upper bound on the student’s performance, particularly when the student’s base model is not more capable than the teacher’s. Using supervised fine-tuning, LN-Ultra can approach the performance of DeepSeek-R1 but not exceed it. To enable students to surpass their teachers, large-scale reinforcement learning is a viable approach, as it allows the model to continually explore new possibilities and engage in self-learning. Consistent with the findings of DeepSeek-AI et al. (2025), our preliminary experiments indicate that applying RL to smaller models yields suboptimal results compared to distillation.Due to resource constraints, we apply reasoning RL only to LN-Ultra, which results in a model that outperforms its teacher.
@@ 7.1 Benchmarks
All evaluations are conducted at a 32k context length, even though training was performed with a maximum sequence length of 16k for LN-Super and 24k for LN-Ultra. We observed consistent improvements when evaluating at expanded context lengths, as shorter sequence limits can truncate long reasoning traces and lead to incomplete generations—particularly on benchmarks that require multi-step reasoning. We use temperature 0.6 and top-p 0.95 for reasoning-on evaluations, and temperature 0 (greedy decoding) for reasoning-off. We generate up to 16 completions per prompt and report average pass@1 accuracy. Checkpoints are selected based on performance on a subset of reasoning benchmarks. As observed in prior works  (Moshkov et al., 2025), evaluation on reasoning-heavy tasks such as AIME can exhibit high variance due to small dataset size and generation randomness. Reported numbers may vary across repeated runs or sampling strategies.
@@ 7.2 LN-Nano
Table 3 shows that LN-Nano achieves strong performance across all reasoning benchmarks, including AIME25-I and LiveCodeBench, despite its small size. This demonstrates the effectiveness of our SFT pipeline and curated reasoning datasets in transferring structured reasoning to compact models. For Nano reasoning SFT blends, carefully balancing data-distribution across math, coding, and stem areas has been important to achieve near state-of-the-art accuracies at SFT stage. For example, our early experiments showed worse accuracies especially in Chemistry related questions, i.e. one of the major areas in GPQA-D. Upsampling Chemistry related data samples in the STEM subset of the overall SFT blend helped to achieve higher GPQA-D accuracies. The RPO stages at the end of the post-training pipeline mainly targeted IFEval accuracy improvement as shown in Table 3.
@@ 7.3 LN-Super
Table 4 compares LN-Super to other models in its weight class. It performs competitively across both reasoning and non-reasoning tasks. In reasoning-off mode, LN-Super performs on par with Llama-3.3-70B, the model it was originally distilled from. In reasoning-on mode, it outperforms competing models such as DeepSeek-R1-Distilled-Llama-70B, providing strong reasoning capabilities without sacrificing instruction following. These results show that this single model offers the strengths of both reasoning-optimized and non-reasoning models, making it effective for general assistant and structured reasoning use cases. Additionally, as shown in Table 4, reasoning-focused SFT causes a noticeable drop in IFEval scores. To recover instruction-following capabilities, we apply a dedicated IFEval RL run (see Section 6.1) to ensure that strong reasoning does not come at the cost of degraded general assistant behavior. Our experimental results reveal another trade-off: optimizing for instruction following (as measured by IFEval) can compromise conversationality (as measured by Arena-Hard), and conversely, prioritizing conversationality may detract from instruction following performance. To address this, we applied model merging to LN-Super, selecting a checkpoint from the Pareto frontier that balances these objectives. Due to mixed outcomes, we did not adopt this approach for other models. The only area where LN-Super underperforms is on LiveCodeBench, which is attributable to its SFT phase being conducted on an earlier version of the dataset, unlike LN-Nano and LN-Ultra. We plan to address this and improve coding-related reasoning performance in a future model refresh.
@@ 9 Conclusions
Denys Fridman, Erick Galinkin, Michael Evans, Katherine Luna, Leon Derczynski, Nikki Pope, Eileen Long, Seth Schneider, Guillermo Siman, Tomasz Grzegorzek, Pablo Ribalta, Monika Katariya, Soumye Singhal*

### v2

@@ 2.1 Deployment Constraints and Efficiency Targets
Figure 4 illustrates the trade-off between GPQA-Diamond accuracy (%) and processing throughput (tokens/s) for two settings. Notably, LN-Ultra consistently outperforms DeepSeek-R1 and Llama-3.1-405B in both accuracy and efficiency across these settings, clearly positioning it as a superior choice on the accuracy-throughput Pareto curve.
@@ 4.2 Model-Specific Training
differently from other models below, undergoes a three-stage SFT pipeline using a global batch size of 256 using sequence packing with effective sequence length of 32k tokens. In the first stage, the model is fine-tuned exclusively on reasoning data from code, math, and science domains (Section 3.1) with a learning rate of $1\mathrm{e}{-4}$ for four epochs. This prevents failure modes such as repetitive completions. In the second stage, we introduce non-reasoning data (Section 3.2) mixed with reasoning samples, allowing the model to learn reasoning control. In the final stage, a smaller blend focused on chat, instruction-following, and tool-calling is used.
@@
is trained on the full SFT dataset for a single epoch using a fixed learning rate of $5\mathrm{e}{-6}$, sequence length of 16k and a global batch size of 256. Smaller-scale runs suggested that performance improves up to 3–4 epochs with larger learning rates ($5\mathrm{e}{-5}$), but training was constrained by computational and time limits. Recent works (Wen et al., 2025) show that rejection fine-tuning can further improve performance; however, it does not yield gains in our experiments and is therefore omitted.
@@
is trained on the full dataset using sequence packing with effective sequence length of 24k and a global batch size of 256 to maximize token throughput—an essential strategy when fine-tuning large models with long-context reasoning data. Initial ablation runs indicated that higher learning rates such as $5\mathrm{e}{-5}$ generally improve outcomes, but consistently high learning rates caused training instability, including gradient explosions. To mitigate this, we implement a linear warmup to $1\mathrm{e}{-5}$, followed by cosine decay to $1\mathrm{e}{-6}$ with a warmup ratio of 10%. Despite these measures, training encountered gradient explosions and numerical instability after the first epoch. This required training resumption with reinitialized optimizer states, after which successful convergence was achieved.
@@ 5 RL for Reasoning
As described in Section 4 and illustrated in Table 5, models can acquire strong capabilities through supervised fine-tuning by distilling knowledge from powerful teachers. However, distillation inherently sets an upper bound on the student’s performance, particularly when the student’s base model is more capable than the teacher’s. Using supervised fine-tuning, LN-Ultra can approach the performance of DeepSeek-R1 but not exceed it. To enable student to surpass their teacher, large-scale reinforcement learning is an essential approach, as it allows the model to continually explore new possibilities and engage in self-learning. Consistent with the findings of DeepSeek-AI et al. (2025), our preliminary experiments indicate that reinforcement learning (RL) yields suboptimal results for smaller models compared to distillation. Given these observations and resource constraints, we apply reasoning RL exclusively to LN-Ultra, enabling it to surpass its teacher model and setting a new state-of-the-art on GPQA amoung open models.
@@ 7.1 Benchmarks
All evaluations are conducted at 32k context length, even though training was performed with a maximum context length of 16k for LN-Super and 24k for LN-Ultra. We observed consistent improvements when evaluating at expanded context lengths, as shorter limits can truncate long reasoning traces and lead to incomplete generations—particularly on benchmarks that require multi-step reasoning. We use temperature 0.6 and top-p 0.95 for reasoning-on evaluations, and temperature 0 (greedy decoding) for reasoning-off. We generate up to 16 completions per prompt and report average pass@1 accuracy. Checkpoints are selected based on performance on a subset of reasoning benchmarks. As observed in prior works  (Moshkov et al., 2025), evaluation on reasoning-heavy tasks such as AIME can exhibit high variance due to small dataset size and generation randomness. Reported numbers may vary across repeated runs or sampling strategies.
@@ 7.2 LN-Nano
Table 3 shows that LN-Nano achieves strong performance across all reasoning benchmarks, including AIME25-I and LiveCodeBench, despite its small size. This demonstrates the effectiveness of our SFT pipeline and curated reasoning datasets in transferring structured reasoning to compact models. For LN-Nano, carefully balancing data-distribution across math, coding, and stem areas has been important to achieve near state-of-the-art accuracies at the SFT stage. For example, our early experiments showed worse accuracies especially in chemistry related questions, one of the major areas in GPQA-D. Upsampling chemistry related data samples in the STEM subset of the overall SFT blend helped to achieve higher GPQA-D accuracies. The RPO stages at the end of the post-training pipeline mainly targeted IFEval accuracy improvement as shown in Table 3.
@@ 7.3 LN-Super
Table 4 compares LN-Super to other models in its weight class where it performs competitively across both reasoning and non-reasoning tasks. In reasoning-off mode, LN-Super performs on par with Llama-3.3-70B, the model it based on. In reasoning-on mode, it outperforms competing models such as DeepSeek-R1-Distilled-Llama-70B, providing strong reasoning capabilities without sacrificing instruction following. These results show that this single model offers the strengths of both reasoning-optimized and non-reasoning models, making it effective for general assistant and structured reasoning use cases. Additionally, as shown in Table 4, reasoning-focused SFT causes a noticeable drop in IFEval scores. To recover instruction-following capabilities, we apply a dedicated IFEval RL run (see Section 6.1) to ensure that strong reasoning does not come at the cost of degraded general assistant behavior. Our experimental results reveal another trade-off: optimizing for instruction following (as measured by IFEval) can compromise conversationality (as measured by Arena-Hard), and conversely, prioritizing conversationality may detract from instruction following performance. To address this, we applied model merging to LN-Super, selecting a checkpoint from the Pareto frontier that balances these objectives. Due to mixed outcomes, we did not adopt this approach for other models. The only area where LN-Super underperforms is on LiveCodeBench, which is attributable to its SFT phase being conducted on an earlier version of the dataset, unlike LN-Nano and LN-Ultra. We plan to address this and improve coding-related reasoning performance in a future model refresh.
@@ 9 Conclusions
Soumye Singhal*, Denys Fridman, Erick Galinkin, Michael Evans, Katherine Luna, Leon Derczynski, Nikki Pope, Eileen Long, Seth Schneider, Guillermo Siman, Tomasz Grzegorzek, Pablo Ribalta, Monika Katariya, Chris Alexiuk